OpenStudy (anonymous):

The equation of a line is . What is the equation of the line that is parallel to the first line and passes through (4, –2)?

OpenStudy (anonymous):

So what's the equation for first line?

OpenStudy (anonymous):

y=1/4x-2

OpenStudy (anonymous):

so slope of first line is the coefficient of x i.e. 1/4 and as it is parellal to the required line. slopes would be same. required eqn is, (y-y1)=m(x-x1) where m=slope and(x1,y1) is the given point (y-(-2))=1/4(x-4) 4(y+2)=x-4 4y+8=x-4 4y-x=-4-8 4y-x=-12